OpenStudy (anonymous):

solve by substitution y=2x-10

OpenStudy (p0sitr0n):

to have a substitution, we need 2 equations

OpenStudy (anonymous):

oh

OpenStudy (p0sitr0n):

do you have 2 equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=2x-10 2y=x-8

OpenStudy (anonymous):

hello

OpenStudy (p0sitr0n):

ah kk nice so replace y into the second equation 4x-20=x-8 3x=12 x=4

OpenStudy (anonymous):

thx!

OpenStudy (p0sitr0n):

no prob